# Fluctuations in random field Ising models

### v1

@@ 2.4 Random field Ising models
Recall the RFIM introduced in 1.1. In this section we assume that the base measures $\mu_{i}=\mu$ for all $i\in[n]$, where the common measure $\mu$ is supported on $[-1,1]$. Consequently, we omit the subscript $i$ in the $\psi_{i}$s (see 1.1). We also assume that $\mathbf{c}=(c_{1},\ldots,c_{n})$ are i.i.d. random variables from a distribution $F$, where $F$ is a distribution on reals. Throughout this section we assume that $\operatorname{\mathbb{E}}_{F}\psi^{\prime}(c_{1})=0$. In particular this assumption holds if both the measures $\mu$ and $F$ are symmetric. Thus, this can be viewed as a generalization of the constant field Ising model where $F$ is degenerate at 00 (See [13, 18]). In the next theorem, for simplicity of presentation, we assume that there exists an approximate eigenvalue $\lambda$ (free of $n$) and a sequence of approximate delocalized eigenvectors $\mathbf{q}\in\mathbb{R}^{n}$, i.e.
$\lVert\mathbf{q}\rVert=1,\quad\lVert\mathbf{A}_{n}\mathbf{q}-\lambda\mathbf{q}\rVert\to 0,\quad\lVert\mathbf{q}\rVert_{\infty}\to 0.$ (2.16)
We will provide specific choices of $\lambda$ and $\mathbf{q}$ in examples that follow. Recall the definitions of $\alpha_{n}$ and $\psi(\cdot)$ from (1.4) and 1.1 respectively. In our next result, we provide both quenched (conditional on $\mathbf{c}$) and annealed (unconditional) central limit theorems for the linear statistic $T_{n}$ after appropriate centering.
@@
Assume ${\bf A}_{n}$ satisfy the MHT condition (2.2). Set $\upsilon:=\operatorname{\mathbb{E}}_{F}\psi^{\prime\prime}(c_{1})>0$, and define
@@
    Here, the RHS is $o_{p}(1)$ when (2.16) holds and $\alpha_{n}\ll\frac{1}{\sqrt{n}}$.
@@
    $d_{KS}\left(\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}\sigma_{i},W_{\infty}+\widetilde{W}_{\infty}\right)$ $=O\Big{(}\textsc{Err}(\mathbf{A}_{n},\mathbf{q})+\sqrt{n\operatorname{\alpha_{\mathnormal{n}}}}\|\bm{\epsilon}\|\Big{)},$ (2.17)
    where $W_{\infty}$ and $\widetilde{W}_{\infty}\sim\operatorname{N}\left(0,\frac{\operatorname{\mathbb{E}}_{F}[\psi^{\prime}(c_{1})^{2}]}{(1-\lambda\upsilon)^{2}}\right)$ are independent. Here, the RHS is $o_{p}(1)$ and $o(1)$ respectively, under (2.16), $\alpha_{n}\ll\frac{1}{\sqrt{n}}$, and $\lVert\bm{\epsilon}\rVert=o\left(\frac{1}{\sqrt{n\operatorname{\alpha_{\mathnormal{n}}}}}\right)$.
@@
In particular, for the usual Ising model with binary spins, one takes $\mu$ to be supported on ±$\pm 1$ with $\mu(\pm 1)=\frac{1}{2}$, which gives $\psi^{\prime}(c)=\tanh(c)$ and $\psi^{\prime\prime}(c)=\operatorname{sech}^{2}(c)$. The annealed limiting normal distribution in (2.17) (without any centering) might, at first glance, seem unfamiliar given the nature of our main theorems earlier. However, note that, due to the independence of $\{c_{i}\}_{i\in[n]}$, the sum $\sum_{i=1}q_{i}\psi_{i}^{\prime}(c_{i})$ converges to a centered normal distribution by a standard Lindeberg-Feller argument. Combining this observation with 2.5 yields the required annealed limit. This also highlights the benefits of refining the centering in 2.5 relative to 2.4, albeit under slightly stronger assumptions.
@@
To see the tightness of our bounds, take $\mathbf{q}=n^{-1/2}\mathbf{1}$ and $\mathbf{A}_{n}$ to be the adjacency matrix of a complete graph scaled by its degree $n-1$. In other words, consider the sample mean under the Curie-Weiss model. Then, we have $\alpha_{n}\lesssim\frac{1}{n}$. Consequently, it follows from 2.6 that
@@
The $n^{-1/2}$ rate matches the well-known Berry-Esseen rate for the sample mean of iid random variables [17], so our rates are tight in this example. Also, if $\alpha_{n}\asymp n^{-1/2}$, it is known that the above CLT fails for the choice $c_{i}=c$ for all $i\in[n]$ as mentioned in Remark 2.3. In general, the rate gets slower as $\alpha_{n}$ and $\lVert\mathbf{q}\rVert_{\infty}$ increase.
@@
[…]
@@
[…]
@@
An interesting feature of 2.7(b) is that if ${\bm{q}}$ satisfies (2.18) (i.e. it is approximately a contrast), the limit distribution is universal, i.e., it does not depend on $\mathbf{q}$. This is because, for all $\mathbf{q}$ satisfying (2.18), $(\mathbf{q},0)$ forms an approximate eigenvector-eigenvalue pair in the sense of 2.6.
@@
Let $\mathnormal{\bm{\sigma}}$ be an observation from the model (1.1), where the matrix $\mathbf{A}_{n}$ is obtained from a $d(n)$ regular graph as described above. Assume further that $d(n)\gg n^{-1/2}$ and $|\theta|<1$.
@@
We note that (2.19) is an expander type condition, which is satisfied by several regular graph ensembles of common interest, such as random regular graphs and Ramanujan graphs.
@@
Example 2.3 (Irregular/unbalanced complete bipartite graph).
[…]
@@
Let $\mathnormal{\bm{\sigma}}$ be an observation from the model (1.1), where the matrix ${\bf A}_{n}$ is obtained from a bipartite graph as described above. Then for $|\theta|<2\min{(h,1-h)}$ and $t\in\{-1,1\}$, we have:
[…]
In our final example, we give a spin-glass model, where the entries of $\mathbf{A}_{n}$ are both positive and negative. For simplicity, we only consider the case when the base measure $\mu_{i}=\mu$ puts equal mass on ±$\pm 1$.
Example 2.4 (Hopfield model / sample covariance matrix).
Suppose the base measure $\mu_{i}\equiv\mu$ in (1.1) is the Rademacher distribution. Let $\{Z_{k,i}\}_{k\in[N],i\in[n]}$ be independent mean 00 and variance $1$ sub-Gaussian random variables. Define a symmetric matrix ${\bf A}_{n}$ with 00 on the diagonals, by setting
${\bf A}_{n}(i,j):=\frac{1}{N}\sum_{k=1}^{N}Z_{k,i}Z_{k,j}$
for $i\neq j$.
@@
Let $\operatorname{\mathnormal{\bm{\sigma}}}$ be an observation from the model (1.1), where the matrix ${\bf A}_{n}$ is coming from the off-diagonal part of a sample covariance matrix (described as above), with $\mu_{i}$ denoting the symmetric Rademacher distribution. Let $\mathbf{q}$ be any vector such that $\lVert\mathbf{q}\rVert=1$ and $\lVert\mathbf{q}\rVert_{\infty}\to 0$.
If $N\gg n^{3/2}$, then we have:
[…]

### v2

@@ 2.3 Central Limit Theorems
2.4 Proof overview and main challenges
This paper studies concentration and CLT for linear statistics in a Gibbs measure with a quadratic Hamiltonian. The centering for both the concentration and CLT is provided by the vector $\bf u$}, which is the expectation of the Mean-Field variational approximation to this Gibbs measure. Under a suitable high temperature assumption on the interaction matrix of the quadratic statistic, the mean field optimizer is unique. However, as opposed to a Dobrusin type high temperature assumption which involves the $\ell_{\infty}$ operator norm, this paper works with the $\ell_{4}$ operator norm, which is much less restrictive. In particular, this allows us to verify a CLT for a broader class of asymptotic regimes, as demonstrated in Example 2.4. We stress that our techniques allow for both ferromagnetic and anti-ferromagnetic models, as well as some spin glass models, as demonstrated via examples.
Our first main result is a concentration for linear statistics (see 2.3), centered via the Mean-Field centering. The starting point of our proof is a concentration for linear statistics with a centering via conditional expectation
$\operatorname{\mathbb{E}}[\sigma_{i}|(\sigma_{j},j\neq i)]=\psi^{\prime}(m_{i}+c_{i}),$
which we derive using the method of exchangeable pairs. Here $m_{i}=\sum_{j=1}^{n}\mathbf{A}_{n}(i,j)\sigma_{j}$ is the local field at $i$, and we assume common base measures $\mu_{i}=\mu$ to simplify the notation $\psi_{i}=\psi$. The main challenge is to derive a concentration with respect to the Mean-Field centering, given by an implicit vector $\bf u$} which satisfies the fixed point equation
$u_{i}=\psi^{\prime}(s_{i}+c_{i}),\qquad\text{ where }\qquad s_{i}=\sum_{j=1}^{n}\mathbf{A}_{n}(i,j)u_{j}.$
In order to get a concentration with respect to $\bf u$} we can write
$\sum_{i=1}^{n}q_{i}(\sigma_{i}-u_{i})=\sum_{i=1}^{n}q_{i}(\sigma_{i}-\psi^{\prime}(m_{i}+c_{i}))+\sum_{i=1}^{n}q_{i}(\psi^{\prime}(m_{i}+c_{i})-u_{i}).$
The first term in the RHS above can be controlled using the conditionally centered concentration. Proceeding to bound the second term, a Taylor’s series expansion gives
$\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}($ $\psi^{\prime}(m_{i}+c_{i})-u_{i})=\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}(\psi^{\prime}(m_{i}+c_{i})-\psi^{\prime}(s_{i}+c_{i}))$
$=\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}(m_{i}-s_{i})\psi^{\prime\prime}(s_{i}+c_{i})+\frac{1}{2}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}(m_{i}-s_{i})^{2}\psi^{\prime\prime\prime}(\xi_{i}+c_{i})$
$=\sum_{\mathnormal{i,j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{i}(\sigma_{j}-u_{j})\psi^{\prime\prime}(s_{i}+c_{i})+O\Big(\sum_{\mathnormal{i}=1}^{\mathnormal{n}}|q_{i}|(m_{i}-s_{i})^{2}\Big)$
$=\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\tilde{q}_{j}(\sigma_{j}-u_{j})+O\Big(\sqrt{\sum_{\mathnormal{i}=1}^{\mathnormal{n}}(m_{i}-s_{i})^{4}}\Big),$
where
$\tilde{q}_{j}:=\sum_{i=1}^{n}\mathbf{C}_{n}(i,j)q_{i},\qquad\text{ with }\qquad\mathbf{C}_{n}(i,j)=\mathbf{A}_{n}(i,j)\psi^{\prime\prime}(s_{i}+c_{i}).$
Thus we have reduced the problem of concentration of ${\bf q}^{\top}\operatorname{\mathnormal{\bm{\sigma}}}$ to that of ${\bf q}^{\top}\mathbf{C}_{n}\operatorname{\mathnormal{\bm{\sigma}}}$. In this way, a recursive argument allows us to focus on the behavior of ${\bf q}^{\top}\mathbf{C}_{n}^{\ell}\operatorname{\mathnormal{\bm{\sigma}}}$ for some large positive integer $\ell$. Under the high temperature assumption we have that above sequence of random variables converge to 0 geometrically fast, thereby allowing us to derive the desired concentration about $\mathbf{u}$. Controlling the error term $\sum_{i=1}^{n}(m_{i}-s_{i})^{4}$ necessitates the high temperature assumption $\|\mathbf{A}_{n}\|_{4}\leq\rho<1$ via the fourth operator norm. We note here that if we had additional structure on $\bf q$} (say $\|{\bf q}\|_{\infty}\lesssim n^{-1/2})$, then the error bound could have been controlled by the weaker high temperature assumption $\|\mathbf{A}_{n}\|\leq\rho<1$. Instead we focus to avoid extra assumptions on the eigenvector, and work with the fourth operator norm.
Next, the proof 2.4 is based on Stein’s method for exchangeable pairs. For simplicity, we sketch the proof assuming that $\mathbf{A}_{n}\mathbf{q}=\lambda_{n}\mathbf{q}$. To elaborate, we first construct an exchangable pair $(\operatorname{\mathnormal{\bm{\sigma}}},\operatorname{\mathnormal{\bm{\sigma}}}^{\prime})$ by the usual way:
Let $I$ be a randomly sampled index from $\{1,2,\ldots,n\}$. Given $I=i$, replace $\sigma_{i}$ with an independently generated $\sigma^{\prime}_{i}$ drawn from the conditional distribution of $\sigma_{i}$ given $\sigma_{j},\ j\neq i$. Then, setting
$T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}}^{\prime})=\mathbf{q}^{\top}(\operatorname{\mathnormal{\bm{\sigma}}}^{\prime}-\mathbf{u}),\quad\text{we have}\quad T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}})-T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}}^{\prime})=q_{I}(\sigma_{I}-\sigma_{I}^{\prime}).$
Noting that $\operatorname{\mathbb{E}}[\sigma_{i}^{\prime}\mid\mathnormal{\bm{\sigma}}]=\psi^{\prime}(m_{i}+c_{i})$, a one term Taylor’s series gives
$\operatorname{\mathbb{E}}(T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}})-T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}}^{\prime})\mid\operatorname{\mathnormal{\bm{\sigma}}})$ $=\frac{1}{n}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}(\sigma_{i}-\psi^{\prime}(m_{i}+c_{i}))$
$\approx\frac{1}{n}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}\left(\sigma_{i}-u_{i}-(m_{i}-s_{i})\psi^{\prime\prime}(s_{i}+c_{i})\right)$
$=$ $\frac{T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}})}{n}-\frac{1}{n}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}\psi^{\prime\prime}(s_{i}+c_{i})\Big(\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)(\sigma_{j}-u_{j})\Big)$
which is a good approximation, if $m_{i}\approx s_{i}$ (this is argued in Lemma 3.2). Provided $s_{i}\approx 0$ (which we argue in Lemma 3.3), the RHS above can be approximated as
$\frac{T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}})}{n}-\frac{1}{n}\sum_{\mathnormal{j}=1}^{\mathnormal{n}}\Big(\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{i}\psi^{\prime\prime}(c_{i})\Big)(\sigma_{j}-u_{j})\approx\frac{(1-\lambda_{n}\upsilon_{n})T_{n}^{*}(\operatorname{\mathnormal{\bm{\sigma}}})}{n}.$
Here, the last line follows on noting that for $(c_{1},\cdots,c_{n})$ i.i.d., we have the concentration
$\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{i}\psi^{\prime\prime}(c_{i})\approx\sum_{\mathnormal{i}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)q_{i}\operatorname{\mathbb{E}}\psi^{\prime\prime}(c_{1})=\lambda_{n}q_{j}\operatorname{\mathbb{E}}\psi^{\prime\prime}(c_{1}),$ (2.16)
where the last equality uses the eigenpair assumption $\mathbf{A}_{n}\mathbf{q}=\lambda_{n}\mathbf{q}$. As a proxy for $\operatorname{\mathbb{E}}\psi^{\prime\prime}(c_{1})$, we use the quantity $\upsilon_{n}=\sum_{i=1}^{n}q_{i}^{2}\psi^{\prime\prime}(c_{i})$, which turns out to be more efficient for controlling associated error terms. Now, the above implies that the Stein’s equation holds approximately. The final KS bound in the RHS of (2.13) results from the various approximation errors in the above expansion.
Finally, 2.5, which gives a more tractable form for the centering, follows from Taylor expansion of the Mean-Field optimizers $\mathbf{u}$ and simplifying the expression for $\mathbf{q}^{\top}\mathbf{u}$ using the approximate eigenpair condition. To begin, use a one term Taylor expansion to get
$u_{i}=\psi^{\prime}(s_{i}+c_{i})\approx\psi^{\prime}(c_{i})+s_{i}\psi^{\prime\prime}(c_{i}),$
which gives $\mathbf{q}^{\top}\mathbf{u}$ as
$\mathbf{q}^{\top}\mathbf{u}$ $\approx\mathbf{q}^{\top}\Psi^{\prime}(\mathbf{c})+\sum_{\mathnormal{i}=1}^{\mathnormal{n}}s_{i}q_{i}\psi^{\prime\prime}(c_{i})=\mathbf{q}^{\top}\Psi^{\prime}(\mathbf{c})+\sum_{\mathnormal{i,j}=1}^{\mathnormal{n}}\mathbf{A}_{n}(i,j)u_{j}q_{i}\psi^{\prime\prime}(c_{i})\approx\mathbf{q}^{\top}\Psi^{\prime}(\mathbf{c})+\upsilon_{n}\lambda_{n}\sum_{\mathnormal{j}=1}^{\mathnormal{n}}u_{j}q_{j},$
where the last line above again uses concentration followed by the eigenpair assumption of (2.16). Now, the claim $\mathbf{q}^{\top}\mathbf{u}\approx\frac{\mathbf{q}^{\top}\Psi^{\prime}(\mathbf{c})}{1-\lambda_{n}\upsilon_{n}}$ follows by adjusting the terms.
Even though our main application is the random field Ising model with i.i.d. magnetic fields, our main results are stated for general inhomogeneous fields, deterministic or random. Allowing this broad generality comes at a cost, as now it is difficult to pinpoint the geometry of the space of Mean-Field optimizers $\mathbf{u}$. The MHT assumption comes to the rescue, forcing the uniqueness of optimizer, and also giving concentration for various statistics, which are helpful to verify the CLT, but also of possible independent interest. We believe that using the tools developed in this paper, one can study concentration and CLT of linear statistics under a broader class off magnetic fields $\bf c$} (such as dependent coordinates, which comes up while analyzing the posterior of a high dimensional linear regression model).
@@ 2.5 Random field Ising models
Recall the RFIM introduced in 1.1. In this section we assume that the base measures $\mu_{i}=\mu$ for all $i\in[n]$, where the common measure $\mu$ is supported on [1,1][-1,1]. Consequently, we omit the subscript $i$ in the $\psi_{i}$s (see 1.1). We also assume that $\mathbf{c}=(c_{1},\ldots,c_{n})^{\top}$ are i.i.d. random variables from a distribution $F$, where $F$ is a distribution on reals. Throughout this section let $\operatorname{\mathbb{E}}_{F}$ denote the expectation with respect to the random field $\mathbf{c}$, and assume that $\operatorname{\mathbb{E}}_{F}\psi^{\prime}(c_{1})=0$. In particular this assumption holds if both the measures $\mu$ and $F$ are symmetric. Thus, this can be viewed as a generalization of the constant field Ising model where $F$ is degenerate at 0 (see [19, 28]). Recall the definitions of $\alpha_{n}$ and $\psi(\cdot)$ from (1.4) and 1.1 respectively. In our next result, we provide both quenched (conditional on $\mathbf{c}$) and annealed (unconditional) central limit theorems for the linear statistic $T_{n}$ after appropriate centering.
@@
Assume ${\bf A}_{n}$ satisfies the MHT condition (2.2). Set $\upsilon:=\operatorname{\mathbb{E}}_{F}\psi^{\prime\prime}(c_{1})>0$, and define
@@
[…]
    where $W_{\infty}$ and $\widetilde{W}_{\infty}\sim\operatorname{N}\left(0,\frac{\operatorname{\mathbb{E}}_{F}[\psi^{\prime}(c_{1})^{2}]}{(1-\lambda\upsilon)^{2}}\right)$ are independent.
@@
In particular, for the usual Ising model with binary spins, one takes $\mu$ to be supported on $\pm 1$ with $\mu(\pm 1)=\frac{1}{2}$, which gives $\psi^{\prime}(c)=\tanh(c)$ and $\psi^{\prime\prime}(c)=\operatorname{sech}^{2}(c)$. The annealed limiting normal distribution in (2.17) (without any centering) might, at first glance, seem unfamiliar given the nature of our main theorems earlier. However, note that, due to the independence of $\{c_{i}\}_{i\in[n]}$, the sum $\sum_{i=1}q_{i}\psi_{i}^{\prime}(c_{i})$ converges to a centered normal distribution by a standard Lindeberg-Feller argument. Combining this observation with 2.5 yields the required annealed limit. This also highlights the benefits of refining the centering in 2.5 relative to 2.4, albeit under slightly stronger assumptions.
Remark 2.4 (Eigenvalue-eigenvector).
Continuing from Remarks 2.1 and 2.2, we need an approximate eigenvalue $\lambda$ (free of $n$) and a sequence of approximately delocalized eigenvectors $\mathbf{q}\in\mathbb{R}^{n}$, i.e.
$\lVert\mathbf{q}\rVert=1,\quad\lVert\mathbf{A}_{n}\mathbf{q}-\lambda\mathbf{q}\rVert\to 0,\quad\lVert\mathbf{q}\rVert_{\infty}\to 0$ (2.18)
to get a CLT (i.e., for the bounds in 2.6 to converge to 0). In part (a), the RHS is $o_{P}(1)$ when (2.18) holds and $\alpha_{n}\ll\frac{1}{\sqrt{n}}$. Similarly, in part (b), the RHS is $o_{P}(1)$ and $o(1)$ respectively, under (2.18), $\alpha_{n}\ll\frac{1}{\sqrt{n}}$, and $\lVert\bm{\epsilon}\rVert=o\left(\frac{1}{\sqrt{n\operatorname{\alpha_{\mathnormal{n}}}}}\right)$. We will provide specific choices of $\lambda$ and $\mathbf{q}$ that satisfy (2.18) in the examples that follow.
@@
To see the tightness of our bounds, take $\mathbf{q}=n^{-1/2}\mathbf{1}$ and $\mathbf{A}_{n}$ to be the adjacency matrix of a complete graph scaled by its degree n1n-1. In other words, consider the sample mean under the Curie-Weiss model. Then, we have $\alpha_{n}\lesssim\frac{1}{n}$. Consequently, it follows from 2.6 that
@@
The $n^{-1/2}$ rate matches the well-known Berry-Esseen rate for the sample mean of i.i.d. random variables [27], so our rates are tight in this example. Also, if $\alpha_{n}\asymp n^{-1/2}$, it is known that the above CLT fails for the choice $c_{i}=c$ for all $i\in[n]$ as mentioned in Remark 2.3. In general, the rate gets slower as $\alpha_{n}$ and $\lVert\mathbf{q}\rVert_{\infty}$ increase.
@@
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}(\sigma_{i}-u_{i})\mid\mathbf{c}$ $\overset{d}{\longrightarrow}$ $\operatorname{N}\left(0,\frac{\operatorname{\mathbb{E}}_{F}\psi^{\prime\prime}(c_{1})}{1-\theta\operatorname{\mathbb{E}}_{F}\psi^{\prime\prime}(c_{1})}\right),$
$\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\sigma_{i}$ $\overset{d}{\longrightarrow}$ $\operatorname{N}\left(0,\frac{\operatorname{\mathbb{E}}_{F}\psi^{\prime\prime}(c_{1})}{1-\theta\operatorname{\mathbb{E}}_{F}\psi^{\prime\prime}(c_{1})}+\frac{\operatorname{\mathbb{E}}_{F}[\psi^{\prime}(c_{1})^{2}]}{(1-\theta\operatorname{\mathbb{E}}_{F}\psi^{\prime\prime}(c_{1}))^{2}}\right),$
@@
$\sum_{i=1}^{n}q_{i}(\sigma_{i}-u_{i})\mid\mathbf{c}$ $\overset{d}{\longrightarrow}$ $\operatorname{N}\left(0,\operatorname{\mathbb{E}}_{F}\psi^{\prime\prime}(c_{1})\right),$
$\sum_{i=1}^{n}q_{i}\sigma_{i}$ $\overset{d}{\longrightarrow}$ $\operatorname{N}\left(0,\operatorname{\mathbb{E}}_{F}\psi^{\prime\prime}(c_{1})+\operatorname{\mathbb{E}}_{F}[\psi^{\prime}(c_{1})^{2}]\right).$
@@
An interesting feature of 2.7(b) is that if $\bm{q}$} satisfies (2.19) (i.e. it is approximately a contrast), the limit distribution is universal, i.e., it does not depend on $\mathbf{q}$. This is because, for all $\mathbf{q}$ satisfying (2.19), $(\mathbf{q},0)$ forms an approximate eigenvector-eigenvalue pair in the sense of 2.6. We will now use this observation to show joint convergence for multiple contrasts.
Definition 2.1.
For two functions $h_{1},h_{2}\in L^{2}\big((0,1]\big)$, define their inner product as $\langle h_{1},h_{2}\rangle:=\int_{0}^{1}h_{1}(x)h_{2}(x)dx$.
For any vector ${\bf x}=(x_{1},\cdots,x_{n})$, define a piecewise constant function $g_{\bf x}:(0,1]\mapsto\mathbb{R}$ given by
$g_{\bf x}(t)=x_{i}\qquad\text{ if }\qquad\frac{i-1}{n}<t\leq\frac{i}{n},1\leq i\leq n.$
Corollary 2.8 (Multivariate convergence for contrasts).
Suppose we are in the set up of 2.7. Let
$\mathcal{Q}\in L^{2}\big((0,1]\big)\quad\text{s.t.}\quad\int_{0}^{1}Q(x)dx=0,\quad\int_{0}^{1}Q(x)^{2}dx=1\Big\}$
be the collection of contrast functions. Let $S_{\infty},\tilde{S}_{\infty}:\mathcal{Q}\to\mathbb{R}$ be Gaussian processes with mean 0 and covariance functions
$K(Q_{1},Q_{2}):=[\operatorname{\mathbb{E}}_{F}\psi^{\prime\prime}(c_{1})]\langle Q_{1},Q_{2}\rangle,\quad\tilde{K}(Q_{1},Q_{2}):=\Big[\operatorname{\mathbb{E}}_{F}\psi^{\prime\prime}(c_{1})+[\operatorname{\mathbb{E}}_{F}\psi^{\prime}(c_{1})]^{2}\Big]\langle Q_{1},Q_{2}\rangle,$
respectively.
  1. (a)
    Then, conditional on $\mathbf{c}$, the finite dimensional marginals of the process $\Big\{\sqrt{n}\langle Q,g_{\operatorname{\mathnormal{\bm{\sigma}}}}-g_{\bf u}\rangle,Q\in\mathcal{Q}\Big\}$ converge to the finite dimensional marginals of the Gaussian process $\{S_{\infty}(Q),Q\in\mathcal{Q}\}$.
  2. (b)
    Unconditionally, the finite dimensional marginals of the process $\{\sqrt{n}\langle Q,g_{\operatorname{\mathnormal{\bm{\sigma}}}}\rangle,Q\in\mathcal{Q}\}$ converge to finite dimensional marginals of the Gaussian process $\{\tilde{S}_{\infty}(Q),Q\in\mathcal{Q}\}$.
@@
Let $\mathnormal{\bm{\sigma}}$ be an observation from the model (1.1), where the matrix $\mathbf{A}_{n}$ is obtained from a $d(n)$-regular graph as described above. Assume further that $d(n)\gg n^{1/2}$ and $|\theta|<1$.
@@
We note that (2.20) is an expander type condition, which is satisfied by several regular graph ensembles of common interest, such as random regular graphs and Ramanujan graphs. Also, by repeating the argument in 2.8, one can get joint convergence of multiple contrast vectors.
@@
Example 2.3 ($f$-random graphs [8, 7]).
Let $f:[0,1]^{2}\to[0,1]$ be a symmetric bounded measurable function, and let $Q:[0,1]\to\mathbb{R}$ be a normalized eigenfunction with eigenvalue $\lambda$, i.e.
$\int_{0}^{1}Q(y)^{2}dy=1,\qquad\int_{0}^{1}f(x,y)Q(y)dy=\lambda Q(x),\quad\forall x\in[0,1].$
For $\mathbf{U}:=(U_{1},\ldots,U_{n})^{\top}\stackrel{{\scriptstyle i.i.d}}{{\sim}}\textnormal{Unif}[0,1]$ and $0\leq\gamma<1/2$, consider an adjacency matrix
$\mathbf{G}_{n}(i,j)\stackrel{{\scriptstyle ind}}{{\sim}}\textnormal{Ber}\left(\frac{f(U_{i},U_{j})}{n^{\gamma}}\right),\quad\forall i<j,$
and define $\mathbf{A}_{n}:=\frac{\theta\mathbf{G}_{n}}{n^{1-\gamma}}$. Note that the random graph $\mathbf{G}_{n}$ is dense if $\gamma=0$, and the above scaling allows sparser graphs as well. Also, note that $f$ is not necessarily continuous, thus allowing graphs with block structures such as stochastic block models or complete bipartite graphs. The following corollary shows that our CLTs still hold for such cases.
@@
Let $\mathnormal{\bm{\sigma}}$ be an observation from the model (1.1), where the matrix $\mathbf{A}_{n}$ is the $f$-random graph from the above example. Assume further that $\operatorname{\mathbb{E}}f(U_{1},U_{2})>0$ and $|\theta|<1$. Then, the following holds:
$\frac{1}{\sqrt{n}}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}Q(U_{i})(\sigma_{i}-u_{i})\mid\mathbf{c}$ $\operatorname*{\xrightarrow{\mathnormal{d}}}\operatorname{N}\Big(0,\frac{\operatorname{\mathbb{E}}_{F}\psi^{\prime\prime}(c_{1})}{1-\theta\lambda\operatorname{\mathbb{E}}_{F}\psi^{\prime\prime}(c_{1})}\Big)$
$\frac{1}{\sqrt{n}}\sum_{\mathnormal{i}=1}^{\mathnormal{n}}Q(U_{i})\sigma_{i}$ $\operatorname*{\xrightarrow{\mathnormal{d}}}\operatorname{N}\Big(0,\frac{\operatorname{\mathbb{E}}_{F}\psi^{\prime\prime}(c_{1})}{1-\theta\lambda\operatorname{\mathbb{E}}_{F}\psi^{\prime\prime}(c_{1})}+\frac{\operatorname{\mathbb{E}}_{F}[\psi^{\prime}(c_{1})^{2}]}{(1-\theta\lambda\operatorname{\mathbb{E}}_{F}\psi^{\prime\prime}(c_{1}))^{2}}\Big).$
In our final example, we give a spin-glass model, where the entries of $\mathbf{A}_{n}$ are both positive and negative. For simplicity, we only consider the case when the base measure $\mu_{i}=\mu$ puts equal mass on $\pm 1$.
Example 2.4 (Diluted Hopfield model/covariance matrix).
Suppose the base measure $\mu_{i}\equiv\mu$ in (1.1) is the Rademacher distribution. Let $\{Z_{k,i}\}_{k\in[N],i\in[n]}$ be independent mean 0 and variance $1$ sub-Gaussian random variables. Define a symmetric random matrix ${\bf M}_{n}$ by setting
${\bf M}_{n}(i,j):=\frac{1}{N}\sum_{k=1}^{N}Z_{k,i}Z_{k,j}.$
Let $\mathbf{G}_{n}$ be the $n\times n$ adjacency matrix of a simple graph (with $\mathbf{G}_{n}(i,i)=0$), and define the coupling matrix $\mathbf{A}_{n}$ as the Hadamard product of $\theta\mathbf{M}_{n}$ and $\mathbf{G}_{n}$, for some inverse-temperature parameter $\theta\in\mathbb{R}$:
$\mathbf{A}_{n}(i,j):=\theta\mathbf{M}_{n}(i,j)\mathbf{G}_{n}(i,j).$
This is the so-called diluted Hopfield model [9, 10] that considers the Hopfield model on graphs. Note that taking $\mathbf{G}_{n}$ as the adjacency matrix of a complete graph results in usual Hopfield model without any dilution [31], with $\mathbf{A}_{n}(i,j)=\theta\mathbf{M}_{n}(i,j)$ for $i\neq j$.
@@
Let $\operatorname{\mathnormal{\bm{\sigma}}}$ be an observation from the model (1.1), where the matrix ${\bf A}_{n}$ is coming from the above diluted Hopfield model, with $\mu_{i}$ denoting the symmetric Rademacher distribution. Let $\mathbf{q}$ be any vector such that $\lVert\mathbf{q}\rVert=1$ and $\lVert\mathbf{q}\rVert_{\infty}\to 0$. Under the asymptotic scaling $N\gg n^{3/2}$ and any inverse-temperature parameter $\theta$ (not depending on $n$), we have:
$\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}\left(\sigma_{i}-u_{i}\right)\mid\mathbf{c}\overset{d}{\longrightarrow}\operatorname{N}\left(0,\operatorname{\mathbb{E}}_{F}\operatorname{sech}^{2}(c_{1})\right),\quad\sum_{\mathnormal{i}=1}^{\mathnormal{n}}q_{i}\sigma_{i}\overset{d}{\longrightarrow}\operatorname{N}\left(0,{1}\right).$